# High-resolution direct phase control in the spectral domain in ultrashort pulse lasers for pulse-shaping applications

### v2

@@ V Conclusion and Next Steps
Efforts in the near future will also see a form of this method integrated into the suite of diagnostics on the GALADRIEL system at General Atomics. It will be used for both on-shot diagnostics of the laser pulse as well as part of a control loop for the laser, helping improve stability for tests that involve pulse shaping. As a part of this control loop, it will eventually enable system users to specify arbitrary pulse shapes without needing to know the exact spectrum and phase needed to create them.
@@
VI Acknowledgements
S.B., M.M., and A.A. were supported in part by the National Science Foundation–Czech Science Foundation partnership (NSF Grant No. PHY-2206777).

### v3

@@ 6 Conclusion
The implementation of DPC pulse shaping on GALADRIEL is still being developed. It is intended to be used for both on-shot diagnostics of the laser pulse as well as part of a control loop for the laser, helping improve stability for tests that involve pulse shaping. As a part of this control loop, it will eventually enable system users to specify arbitrary pulse shapes without the user needing to know the exact spectrum and phase needed to create them.
@@
7 Acknowledgements
S.B., M.M., and A.A. were supported in part by the National Science Foundation–Czech Science Foundation partnership (NSF Grant No. PHY-2206777). GALADRIEL access provided by General Atomics internal research and development.